The tensions on the tight side and slack side of the flat belt are 750N and 250N

respectively. If the speed of the belt is 2m/s and the coefficient of friction between the

belt and the pulley is 0.3, neglecting centrifugal tension,

Determine; (i) Power transmitted at this speed in kW.

(ii) Angle of lap in degrees


1
Expert's answer
2021-04-16T07:47:39-0400

Power transmitted, P "\frac{T_1-T_2 \times v}{1000}kW"


"Power =\frac{750-250 \times v}{1000}kW"


"Power =1kW"


Angle of lap

"\frac{T_1}{T_2}=e^{\mu \theta}"

"\ln({\frac{T_1}{T_2}})= \mu \theta"

"1.0986 =0.3 \times \theta"

"\theta= \frac{1.0986}{0.3}=3.662"

"\theta =3.662 \times \frac{180}{\pi}=209.78^0"